# Kinetically Consistent Coarse Graining using Kernel-based Extended Dynamic Mode Decomposition

### v1

@@ 4.1.3 Identification of CG dynamics
Having effective drift and diffusion, we are able to build the learned SDE governing the CG coordinate. We use the Euler-Maruyama scheme to integrate the learned and reference SDEs with integration time step of $dt=10^{-3}$. Figure 2 d shows two trajectories of the CG coordinate $\phi$ for both dynamics for $10^{4}$ time steps, using the same Brownian motion for both trajectories. The propagated learned system follows the reference closely, with both systems staying long times in each meta stable state, and rarely swapping in between those. The results demonstrate that the proposed method can approximate the full system’s meta stable sets well, and identify a suitable SDE for CG dynamics which is accurate even on the level of individual trajectories.
@@
In this example, both the full dynamics and the learned effective dynamics possess state-dependent diffusion fields. We compare the properties of the learned CG model with variable diffusion to those of a CG dynamics with constant diffusion. For the latter, we set the effective diffusion constant to $a=\frac{2}{\beta}=2$. We propagate the corresponding SDEs for a sufficiently large span of time, and estimate a new generator EDMD model based on these simulations, this time using Random Fourier Features as observables 44 . Figure 3, shows the eigenvalues of the generator for these cases compared to the learned generator built upon the original dataset. The result shows that learning a state-dependent diffusion is necessary to recover the original system’s leading eigenvalues. To show the scalability of the method, we repeat the analysis on a real system in the following section.
Figure 3: Dominant eigenvalues of the generator, using models built on simulation data of the learned coarse-grained dynamics with state-dependent diffusion (SDD, orange) and with constant diffusion (CD, green). As a comparison, we show the eigenvalues of the reference generator $\hat{\mathbf{L}}_{r}$ using the original dataset (blue triangles). Note that the first eigenvalue is omitted as it is zero.
@@ 4.2.1 System introduction
$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi(x)\frac{2}{\beta\gamma}M^{-1}\nabla_{x}\xi^{\top}(x)$ (47)
where $M$ is the diagonal mass matrix of all atoms, $\beta$ is the inverse temperature, $\gamma$ is the friction, and $\nabla_{x}\xi(x)$ is the Jacobian of dihedral angles. This way of computing the local diffusion corresponds to a reversible overdamped Langevin process in full position space, meaning that the resulting CG model is an approximation to that same overdamped process. Fitting an overdamped model is simpler than modeling the full non-reversible underdamped Langevin process if a non-linear CG map is used. In addition, the underdamped process is expected to be close to an overdamped model in position space after a re-scaling of time. However, this also means that the eigenvalues and timescales of the overdamped model are expected to differ from the ones obtained by directly approximating the underdamped Langevin model, while the meta stable states will remain the same.
The free energy landscape of the system with respect to these two angles is shown in Figure 4.
@@
The free energy landscape shows four minima, two on the left side, and two in the central part. We apply the kgEDMD algorithm to find the meta stable sets, and then use Algorithm 1 to learn the effective force and a state-dependent effective diffusion field in the dihedral angle space. Because of the periodicity of the CG coordinates, $\phi$ and $\psi$, we adopt the periodic Gaussian kernel. Similar to the previous example, we tune the bandwidth in the kernel function using the VAMP-score.
4.2.2 Meta stability analysis
Figure 5 a shows the leading eigenvalues of the generator obtained from the kgEDMD model.
Figure 5: Approximation of generator for alanine dipeptide. The dominant eigenvalues of the reference generator $\hat{\mathbf{L}}_{r}$ and the learned generator $\hat{\mathbf{L}}_{\alpha}^{\xi}$ built upon the learned effective diffusion on the left, and the relative error of these eigenvalues compared to the reference is shown on the right.
The figure indicates the four dominant eigenvalues which are corresponding to the four minima in the free energy landscape followed by a gap. In addition, we also show the eigenvalues of the generator $\mathbf{L}^{\xi}_{\alpha}$ corresponding to the optimal effective diffusion, which agree well with the reference. Errorbars in the figures are generated by analyzing $5$ independent subsampled sets of the original data set, each one comprising $2000$ samples.
4.2.3 Identification of CG dynamics
For this 2-dimensional CG dynamics, we can express the diffusion field in a $2\times 2$ full matrix. Here we assume though, that the learned diffusion is a diagonal matrix. Figure 6 shows the first and second diagonal terms of the learned diffusion field.
Figure 6: The first (a) and second (b) diagonal terms of the learned diffusion covariance matrix $a$, the reference free energy surface (c) and the free energy surface learned via kernel-based force matching (d).
For the diffusion learning process, we incorporate $2000$ samples of the available dataset.
The next step is to apply force matching algorithm to obtain the effective force. The reference and learned effective free energy surface are depicted in Figures 6c and  6d, respectively. It it noticeable that the learned free energy surface correctly captures all energetic minima, and approximates the energy barriers properly. We also note that $m=50,000$ samples were required to obtain the results in Figure 6 d with kernel-based force matching.
@@ 4.2.3 Identification of CG dynamics
Figure 8: Dominant eigenvalues of the generator, using models built on simulation data of the learned coarse-grained dynamics with state-dependent diffusion (Prop.-SDD, green) and with constant diffusion (Prop.-CD, red). As a comparison, we show the eigenvalues of the optimal learned generator matrix $\hat{\mathbf{L}}_{\alpha}^{\xi}$, assuming either state-dependent diffusion (blue dots) or constant diffusion (orange triangles). Note that the first eigenvalue is omitted as it is zero.
@@ 5 Discussion
We presented the use of kernel-based approximation to the Koopman generator – kgEDMD – to determine meta stable sets and to learn the parameters of CG models for stochastic systems. For reversible systems, we presented a single learning pipeline to identify both the potential energy and the diffusion governing the effective SDE in CG space. By means of two examples, a two-dimensional model potential and the alanine dipeptide, we showed that the learned effective dynamics is able to reproduce both thermodynamic and kinetic quantities of the full dynamics with high accuracy.
In the molecular example, we found that the CG dynamics based on the fully state-dependent diffusion performed comparably to a constant-diffusion model using the average value of the learned diffusion. It should be noted, however, that determining this average value requires to learn the optimal state-dependent model in the first place. However, these observations can serve as a basis for a bottom-up construction of the effective diffusion, starting from a simple constant model, and increasing the complexity incrementally. Additional future research topics include improvements to the efficiency of the method, using, for example, sparse approximations using random Fourier features 45 . Moreover, one can also try to simultaneously optimize the CG mapping $\xi$ along with the parameters of the CG model, for instance by balancing the VAMP score versus the complexity of the CG model.
@@ 6.1 Notation
$k,\Phi$ kernel and associated feature map
$\mathbb{H}$ reproducing kernel Hilbert space
@@
$\nabla_{1}f$ gradient of $f$ with respect to the first argument
$\nabla_{2}f$ gradient of $f$ with respect to the second argument
$A\cdot_{|a,b}B$ contraction of arguments $a$-th of $A$ and $b$-th of $B$
@@ 6.2 VAMP-score
We tune hyper-parameters of the proposed method based on VAMP variational principle proposed in 27, stating that for reversible systems, the $k$ dominant eigenvalues of the Koopman generator can be obtained by a minimization problem
@@
where the $\phi_{i}$ are orthogonal functions. We use this variational principle to optimize the kernel bandwidth in the context of our proposed method. To do this robustly and avoid overfitting, we make use of standard cross validation scheme by introducing $40\%$ of dataset as test set. Figure 9 shows the result for the lemon slice example, using Gaussian and periodic Gaussian kernels.
Figure 9: VAMP-score analysis for the Lemon slice example using periodic Gaussian kernel on the left and Gaussian kernel on the right
We applied the same procedure for optimizing the bandwidth for the example of alanine dipeptide. The only difference is that due to the relative complexity of this example compared to the Lemon slice, and requiring to incorporate more data for the optimization, we use random Fourier features (RFF) to reduce the size of basis functions, leading to a low-rank approximation of the generator. Figure 10 shows the result for optimization of the bandwidth.
Figure 10: VAMP-score analysis for alanine dipeptide.

### v2

@@ 4.1.3 Analysis of the CG dynamics
With the effective drift and diffusion fields, we are able to simulate the learned SDE governing the CG coordinate. We use the Euler-Maruyama scheme to integrate the learned and reference SDEs with integration time step of $dt=10^{-3}$. Figure 2d shows two trajectories of the CG coordinate $\phi$ for both dynamics for $10^{4}$ time steps, using the same Brownian motion for both trajectories. The propagated learned system follows the reference closely, with both systems staying long times in each meta-stable state, and rarely swapping in between those. Combined, the results above demonstrate that the proposed method can approximate the full system’s meta-stable sets well, and identify a suitable SDE for CG dynamics which is accurate even on the level of individual trajectories.
@@
As a final analysis, we compare the properties of the learned CG model with variable diffusion to those of a CG dynamics with constant diffusion, in order to demonstrate the necessity of allowing a state-dependent diffusion. We set the effective diffusion for the constant model to $a=\frac{2}{\beta}=2$. We propagate the corresponding SDEs for a sufficiently large span of time, and estimate a new generator EDMD model based on these simulations. Figure 3, shows the eigenvalues of the generator for these cases compared to the learned generator built upon the original dataset. The result shows that learning a state-dependent diffusion is necessary to recover the original system’s leading eigenvalues.
Figure 3: Dominant eigenvalues of the generator, using models built on simulation data of the learned coarse-grained dynamics with state-dependent diffusion (SDD, orange) and with constant diffusion (CD, green). As a comparison, we show the eigenvalues of the generator $\hat{\mathbf{L}}_{r}$ using the original dataset (blue). Note that the first eigenvalue is omitted as it is zero.
@@ 4.2.1 System introduction
Alanine dipeptide is a model system widely used in method development for simulation studies of macro-molecules. Figure 4 shows the graphical representation of Alanine dipeptide. It is well known that the dynamical behavior of the molecule can be expressed in terms of the backbone dihedral angles $\phi$ and $\psi$, which constitute the two-dimensional reaction coordinate space defining the CG map $\xi$:
$\xi(x)=\begin{bmatrix}\phi(x)&\psi(x)\end{bmatrix}.$ (34)
@@
The familiar free energy landscape of the system with respect to these two angles is shown in Figure 4, displaying four minima, two on the left side, and two in the central part.
@@
We apply the gEDMD algorithm with random Fourier features to find the meta-stable sets, and then use Algorithm 1 to learn the effective force and a state-dependent effective diffusion field in the dihedral angle space. Because of the periodicity of the CG coordinates, $\phi$ and $\psi$, the spectral measure corresponds to a periodic Gaussian kernel. Similar to the previous example, we tune the bandwidth of the kernel function using the VAMP-score.
4.2.2 Meta-stability analysis
Figure 5a shows the leading finite timescales by taking reciprocals of the first three nonzero eigenvalues of the generator obtained from the gEDMD matrix $\hat{\mathbf{L}}_{r}$ (error bars in the figure are generated by analyzing $5$ independent subsampled sets of the original data set, each comprising $50000$ samples).
Figure 5: Approximation of generator for alanine dipeptide. The dominant timescales corresponding to the reference generator $\hat{\mathbf{L}}_{r}$ and the learned generator $\hat{\mathbf{L}}_{\alpha}^{\xi}$ built upon the learned effective diffusion on the left, and the relative error of these timescales compared to the reference is shown on the right.
The figure indicates the three dominant timescales which are corresponding to the four minima in the free energy landscape followed by a gap. In addition, we also show the timescales corresponding to the generator $\mathbf{L}^{\xi}_{\alpha}$ based on the optimal effective diffusion, which agree well with the reference. As another way for validation of the timescales, we also computed the leading implied timescales using an MSM trained on the original simulation data, which does not rely on the overdamped assumption. These timescales exceed those of the generator model $\mathbf{L}^{\xi}_{\alpha}$ by a uniform factor of about $100$, their re-scaled values are shown as black dashed lines. Thus, we observe that the dynamics in CG space based on the overdamped assumption is accelerated by a factor $100$ for this example, which confirms similar observations from previous studies. In principle, the exact transition timescales of the full simulation can be recovered by re-scaling the friction term of the CG model accordingly.
4.2.3 Analysis of the CG dynamics
For this 2-dimensional coarse graining, we can express the diffusion field as a $2\times 2$ full matrix. For simplicity, however, we assume that the learned diffusion is a diagonal matrix. Figure 6 shows the first and second diagonal terms of the learned diffusion field based on $50000$ samples of the available dataset.
Figure 6: The first (a) and second (b) diagonal terms of the learned diffusion covariance matrix $a$, the reference free energy surface (c) and the free energy surface learned via KDE (d).
To learn the effective potential, we found that the KDE method works best. The reference and learned effective free energy surfaces are depicted in Figures 6c and  6d, respectively. It it noticeable that the learned free energy surface correctly captures all energetic minima and barriers up to some minor spurious behaviour close to the transition regions. We emphasize once again that this approximation could probably be improved further by using a more accurate learning method.
@@ 4.2.3 Analysis of the CG dynamics
Figure 7: Left: Free energy surface learned via KDE. Right: estimated free surface from histogramming the simulated CG dynamics.
@@
Figure 8: Kinetic consistency of the CG dynamics for alanine dipeptide. Left: PCCA+ membership analysis applied to simulation data of the CG dynamics. Right: slowest finite timescales calculated using an approximation of the generator from the reference dataset (blue) and the propagated CG dynamics with state-dependent diffusion (SDD, orange) as well as constant diffusion (CD, green), compared to those obtained via a Markov state model (black).
4.3 Chignolin
4.3.1 System introduction
@@
$\xi(x)=\begin{bmatrix}\mathrm{TIC}_{1}(x)&\mathrm{TIC}_{2}(x)\end{bmatrix}.$ (35)
By projecting the atomistic positional information of the system onto this $2$-dimensional TICA space and computing the histogram of the data, the free energy surface can be obtained, as shown in Figure 2.
Figure 9: Graphical representation of CLN025 on the left  [^2] The image is generated using Protein Data Bank in Europe platform., and the reference free energy surface in the two-dimensional TICA space on the right. The left-hand side minimum corresponds to the folded state, the bottom right minimum corresponds to the unfolded state and the top one associates to the misfolded state.
The free energy surface shows three minima, representing the three conformational states of folded, unfolded and misfolded. We perform the same analysis as for the previous example.
4.3.2 Meta-stability analysis
To find the timescales of the system, we applied the gEDMD method with random Fourier features as before, and computed the eigenvalues of the generator model $\hat{\mathbf{L}}_{r}$. Figure 10 shows the corresponding timescales of the system, which are the inverse of the generator’s eigenvalues. The figure indicates the two leading timescales of the system (the first one associated to the stationary distribution is not shown) corresponding to the three meta-stable sets, followed by a spectral gap. Moreover, we show that the timescales of the CG generator $\mathbf{L}^{\xi}_{\alpha}$ for the optimal effective diffusion are very similar, the relative errors shown on the right of the same figure are sufficiently small. Also, we observe that the gEDMD timescales are once again uniformly re-scaled compared to the leading timescales of an MSM estimated on the original data, which is not subject to the overdamped assumption. The re-scaling factor is quite drastic this time, reducing micro-second timescales of the full system to less than pico-seconds for the CG dynamics. Nevertheless, as the re-scaling is again uniform, the original timescales can in principle be recovered by re-scaling the friction term. Error-bar figures were again generated by analyzing $5$ independent subsampled sets, each comprising $1.6\times 10^{5}$ samples.
Figure 10: Approximation of generator for Chignolin. The slowest finite timescales corresponding to the reference generator $\hat{\mathbf{L}}_{r}$ and the learned generator $\hat{\mathbf{L}}_{\alpha}^{\xi}$ built upon the learned effective diffusion on the left, and the relative error on the right. The black dashed lines on the left indicate re-scaled timescales obtained from an MSM on the original simulation data.
4.3.3 Analysis of the CG dynamics
Following the same procedure as in the previous examples, we learned a $2\times 2$ diffusion matrix in the CG space, but this time, we tested out a full non-diagonal diffusion field. Figure 11 shows the four elements of the learned diffusion matrix. In addition, the left panel of Figure 12 depicts the free energy surface learned by the KDE method, which is in satisfactory agreement with the reference one in Figure 2.
Figure 11: Components of the learned diffusion covariance matrix $a^{\xi}$ for Chignolin in its two-dimensional TICA space (note that the off-diagonal elements are symmetric).
Figure 12: Free energy surface in the two-dimensional TICA space for Chignolin, as learned by the KDE estimator on the left, and obtained from a histogram of the CG dynamics on the right.
@@ 4.3.3 Analysis of the CG dynamics
As we are mainly interested in kinetic properties, we compute a new gEDMD model on the propagated CG dynamics, and re-compute the associated eigenvalues and eigenvectors. The result of a PCCA+-analysis indicates that the correct meta-stable sets are recovered, as shown in the left panel in Figure 13. Likewise, the leading implied timescales estimated from the simulated CG dynamics are in good agreement with those of the original gEDMD model $\hat{\mathbf{L}}_{r}$ and the re-scaled MSM timescales, both estimated from the original simulation data, as shown in the right panel of Figure 13.
Similar to the previous example, we also generate a separate trajectory based on a constant diffusion according to the average of the learned diffusion. We find that transition timescales for the constant diffusion are not well fitted to the reference. This is likely due to the fact that the variation of the diffusion field is stronger than in the previous example, and taking the mean leads to a slower system. This result confirms the need to learn a state-dependent diffusion field in the CG space to achieve kinetic consistency.
Figure 13: Kinetic consistency of the learned CG model for Chignolin. Left: PCCA+ states obtained from simulating the learned CG model. Right: slowest finite timescales of the system calculated using an approximation of the generator from the reference dataset (blue) and from the propagated CG dynamics (state-dependent diffusion in orange, constant diffusion in green). We also compare to re-scaled timescales from a Markov state model on the original simulation data (black).
@@ 5 Discussion
We presented a novel approach to learn kinetically consistent coarse grained models for stochastic dynamics. We have introduced a learning method for the effective diffusion field in CG space, and shown how the kinetic properties of the CG dynamics can be evaluated by exploiting models for the Koopman generator (gEDMD algorithm). We have also shown that random Fourier features provide an efficient and flexible parametrization for both the effective diffusion and the gEDMD model. By means of three examples, a two-dimensional model potential and two datasets of molecular dynamics simulations, we showed that the effective dynamics in low-dimensional reaction coordinate spaces are able to reproduce both thermodynamic and kinetic quantities of the full dynamics accurately.
For the molecular examples, we have relied on the overdamped assumption to parametrize reversible CG dynamics. We have seen that this assumption leads to a uniform acceleration of the CG dynamics compared to the full system. The re-scaling factor can be estimated numerically by comparing the gEDMD model to a kinetic model that does not rely on the overdamped assumption. We used MSMs in this paper, but note that a more general EDMD model (e.g. using random features) would work just as well.
In this study, we used long equilibrium simulations to train CG models. However, one of the appealing aspects of the generator EDMD approach is that it only requires Boltzmann samples. As has been pointed out in previous studies, these samples can also be obtained from biased sampling simulations 42, 43 , or by employing generative models 44 .
Among other topics, future work will focus on applying the formalism to higher-dimensional and more transferrable CG coordinates, for example C-alpha models. Another topic is the construction of CG models that can explicitly account for the underdamped structure of the full system, or that can incorporate memory terms, which were entirely disregarded in our study. Moreover, one can also try to simultaneously optimize the CG mapping $\xi$ along with the parameters of the CG model, for instance by balancing the VAMP score versus the complexity of the CG model.
Acknowledgments
We thank the Theoretical and Computational Biophysics Group at Freie Universität Berlin for sharing the simulation data of the Chignolin mini-protein.
@@ 6.1 Notation
$A\cdot_{|i,j}B$ contraction of dimensions $i$ and $j$ of arrays $A$ and $B$.
@@ 6.2 VAMP-score
We tune hyper-parameters of the proposed method based on VAMP variational principle proposed in 26 , stating that for reversible systems, the $k$ dominant eigenvalues of the Koopman generator can be obtained by a minimization problem
@@
where the $\phi_{i}$ are orthogonal functions. We use this variational principle to optimize the kernel bandwidth giving rise to the spectral measure used in the context of our proposed method. To do this robustly and avoid overfitting, we make use of standard cross validation scheme by introducing $40\%$ of dataset as test set. Figure 14 shows the result for the lemon slice example, using Gaussian and periodic Gaussian kernels.
Figure 14: VAMP-score analysis for the Lemon slice example using periodic Gaussian kernel on the left and Gaussian kernel on the right.
We applied the same procedure for optimizing the bandwidth for alanine dipeptide and Chignolin. Figure 15 shows the result for optimization of the bandwidth.
Figure 15: VAMP-score analysis for alanine dipeptide on the left and Chignolin on the right.
As the figure indicates, there exist a range of the bandwidth $\gamma$ that we can safely choose the bandwidth from.